OpenStudy (anonymous):

Given the equation y − 3 =1/2 (x + 6) in point-slope form, identify the equation of the same line in standard form. x − 2y = −12 y =1/2 x y = 1/2x + 6 y = 1/2x + 9

OpenStudy (anonymous):

y-3=1/2(x+6) => y-3 = (1/2)x +3 => y=1/2x + 6